# Bochner-type Theorems for Distributional Category

### v1

@@ 4. Extensions for distributional category
In this section, we will see that $\operatorname{{\mathsf{dcat}}}$ allows us to improve the results of Section 3. As motivation, recall that a compact manifold $M^{n}$ is flat if the sectional curvature at each point is zero. It is known that $M^{n}$ is flat if and only if $M\cong{\mathbb{R}}^{n}/G$, where $G$ is a Bieberbach group (see below) acting freely on ${\mathbb{R}}^{n}$ and $\cong$ denotes “is isometric to”. Hence, $M$ is aspherical with fundamental group $\pi_{1}(M)=G$. Bieberbach’s Theorem characterizes a Bieberbach group $G$ as one having a maximal normal free abelian subgroup $N\cong{\mathbb{Z}}^{k}$ such that $G/N=Q$ is finite. In [Va70], Vasquez removed the maximal condition by showing that any free abelian $N$ with $G/N$ finite has a normal maximal free abelian centralizer of the same rank. Furthermore, using the fact that a finite index subgroup contains a normal finite index subgroup (namely the normal core), it can be shown that $G$ is the fundamental group of a compact flat Riemannian manifold of dimension $n$ if and only if
@@
Hence, if $M$ is flat, then $\pi_{1}(M)=G$ obeys these conditions. The converse was shown in [FH83]; that is, a closed (connected) manifold $M^{n}$ ($n\not=3,4$) supports a flat Riemannian structure if and only if $M^{n}$ is aspherical (with $\operatorname{{\rm cd}}(\pi_{1}(M))=n$) and $\pi_{1}(M)$ contains an abelian subgroup of finite index. From Section 2.E, we see that a flat manifold $M$ has the Gottlieb group given by the center of the fundamental group, i.e., $G(M)=\mathcal{Z}\pi_{1}(M)$. Moreover, by Theorems 2.4 and 3.7, we have
@@
In [HS86], it was shown that $\operatorname{{\mathsf{rank}}}(\mathcal{Z}\pi_{1}(M))=b_{1}(\pi_{1}(M))=b_{1}(M)$, so we see that we also have
$b_{1}(M)\leq\operatorname{{\mathsf{dcat}}}(M),$
an analog to Theorem 3.3. In this case, the condition
$b_{1}(M)=\operatorname{{\mathsf{cat}}}(M)=\operatorname{{\mathsf{dcat}}}(M)=\dim(M)=n$
then says that $\pi_{1}(M)$ contains a central free abelian subgroup ${\mathbb{Z}}^{n}$. This can only happen if $\pi_{1}(M)$ is itself free abelian of rank $n$ (essentially by the same argument as in the proof of Theorem 4.2 (3)). Hence, $M=T^{n}$.
Since flat manifolds have non-negative Ricci curvature, they provide non-trivial examples of such manifolds with non-trivial Gottlieb groups, where we can also determine $\operatorname{{\mathsf{dcat}}}$. Also, in [DHS09] many flat manifolds were described that are Kähler (hence symplectic), so we also have non-trivial examples of $c$-symplectic nature too. Finally, we note that many of these flat Kähler manifolds have non-trivial centers (determined by non-zero first Betti numbers), so they have non-trivial Gottlieb groups as well.
@@
Let $M$ be a closed $n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is infinite. If $M^{\prime}=T^{r}\times W$ is a Cheeger–Gromoll splitting of $M$, then
@@
If $M$ is a torus, then $b_{1}(M)=\operatorname{{\mathsf{cat}}}(M)=\operatorname{{\mathsf{dcat}}}(M)$ by Theorems 3.3 and 2.4. Conversely, if $b_{1}(M)=\operatorname{{\mathsf{dcat}}}(M)$, then the above inequalities give $\operatorname{{\mathsf{cup}}}(W)=0$. Since $W$ is closed and simply connected (hence orientable), this implies $W=\ast$. Hence, $M^{\prime}=T^{n}$. This means $M$ is a closed aspherical $n$-manifold. Therefore, $b_{1}(M)=\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)=n$ by Theorem 2.4. But this is the situation of Bochner’s theorem, so we obtain $M=T^{n}$. ∎
@@
Recall that in geometric group theory (see [BB97]), a discrete group $\Gamma$ is said to be in $FH(R)$ (written $\Gamma\in FH(R)$) for a commutative ring $R$ if $\Gamma$ acts freely, properly discontinuously, cellularly, and cocompactly on an $R$-acyclic space (i.e., a space $Y$ with $H^{j}(Y;R)=0$ for $j>0$). This is one step in the ladder of finiteness conditions for a group (also see [Br82]).
@@
Let $p:X^{\prime}\to X$ be a regular covering map corresponding to $A\subset\pi_{1}(X)$, so that $\pi_{1}(X^{\prime})=A$. In fact, by Property (2) in Section 2.E, we have $\pi_{1}(X^{\prime})=G(X^{\prime})$ since $p_{*}(A)\subset G(X)$. Because $A$ is free abelian, we have $A={\mathbb{Z}}^{k}$ for some $k$ so that the Hurewicz rank of $A$ is $k$. Thus, by Theorem 2.5, we have $X^{\prime}\simeq T^{k}\times Y$. Clearly, Theorem 2.2 gives us the following series of inequalities (where recall that $\operatorname{{\mathsf{cup}}}$ denotes the rational cup-length):
@@
Note that in the previous argument, we only needed $X$ to be compact to fit the definition of $FH({\mathbb{Q}})$. The fact that $Y$, and hence $\widetilde{X}$, is ${\mathbb{Q}}$-acyclic does not depend on the compactness of $X$. Now, if $X$ is a nilpotent space (i.e., $\pi_{1}(X)$ is nilpotent and acts nilpotently on the homology of the universal cover), we can rationalize it to get $X_{\mathbb{Q}}$ with the property that $\pi_{1}(X_{\mathbb{Q}})$ is the Malcev ${\mathbb{Q}}$-completion $\pi_{1}(X)_{\mathbb{Q}}$ of the nilpotent group $\pi_{1}(X)$, and $\pi_{j}(X_{\mathbb{Q}})=\pi_{j}(X)\otimes{\mathbb{Q}}=\pi_{j}(\widetilde{X})\otimes{\mathbb{Q}}=0$ for each $j\geq 2$ because $\widetilde{X}$ is simply connected and ${\mathbb{Q}}$-acyclic. Therefore, $X_{\mathbb{Q}}=K(\pi_{1}(X_{\mathbb{Q}}),1)=K(\pi_{1}(X)_{{\mathbb{Q}}},1)$ is an Eilenberg–Mac Lane space and $X$ itself is rationally homotopy equivalent to $K(\pi_{1}(X),1)$. This is often written $X\simeq_{\mathbb{Q}}K(\pi_{1}(X),1)$.
Now, suppose that $X$ is a closed $n$-manifold and $G(X)$ has finite index in $\pi_{1}(X)$. Then we see that $A={\mathbb{Z}}^{k}$ (as above) also has a finite index in $\pi_{1}(X)$. Therefore, $X^{\prime}\simeq T^{k}\times Y\to X$ is a finite covering and $X^{\prime}$ is then a closed $n$-manifold as well. Thus, $Y$ is also a closed manifold of dimension $n-k$. Just as before, since $\operatorname{{\mathsf{rank}}}(G(X))=k=\operatorname{{\mathsf{dcat}}}(X)$, we have $\operatorname{{\mathsf{cup}}}(Y)=0$. Hence, $Y$ is ${\mathbb{Q}}$-acyclic and simply connected. This means that $X^{\prime}$ is an orientable closed $n$-manifold. So, we must have $H_{n}(X^{\prime};{\mathbb{Q}})={\mathbb{Q}}$. But since $Y$ is ${\mathbb{Q}}$-acyclic, this can only happen if $k=n$, $Y\simeq*$, and $X^{\prime}\simeq T^{n}$. This then means that $X\simeq K(\pi_{1}(X),1)$, so that $\pi_{1}(X)$ is torsion-free. But then we have a short exact sequence
@@
where ${\mathbb{Z}}^{n}$ is central in $\pi_{1}(X)$ and $F$ is finite. Consider the associated $5$-term exact sequence (see [Br82, Corollary VII.6.4])
@@
Because ${\mathbb{Z}}^{n}$ is central, the middle term reduces to ${\mathbb{Z}}^{n}$ itself. The quotient $F$ is finite, so $H_{2}(F)$ is finite as well and therefore, the image of $H_{2}(F)\to{\mathbb{Z}}^{n}$ is zero and ${\mathbb{Z}}^{n}\to H_{1}(\pi_{1}(X))$ is an injection. This translates to an injection ${\mathbb{Z}}^{n}\to\pi_{1}(X)/[\pi_{1}(X),\pi_{1}(X)]$ so that ${\mathbb{Z}}^{n}\cap[\pi_{1}(X),\pi_{1}(X)]=\{1\}$. But then we have an injection $[\pi_{1}(X),\pi_{1}(X)]\to\pi_{1}(X)/{\mathbb{Z}}^{n}$, where $\pi_{1}(X)/{\mathbb{Z}}^{n}$ is finite by our hypothesis. Since as a subgroup of $\pi_{1}(X)$, $[\pi_{1}(X),\pi_{1}(X)]$ is torsion-free, we must have $[\pi_{1}(X),\pi_{1}(X)]=\{1\}$. Therefore, $\pi_{1}(X)$ is a finitely generated torsion-free abelian group. This implies that $\pi_{1}(X)={\mathbb{Z}}^{n}$ and $X\simeq T^{n}$. ∎
@@
In [Ja25, Section 3.2], it was shown that $\operatorname{{\mathsf{dcat}}}(\mathcal{KT})=4$ where $\mathcal{KT}=\mathcal{H}_{3}\times S^{1}$ is the Kodaira–Thurston symplectic non-Kähler manifold (and $\mathcal{H}_{3}$ is the $3$-dimensional Heisenberg nilmanifold given by upper-triangular $3\times 3$-matrices in $GL({\mathbb{R}},3)$ with $1$’s on the diagonal modulo such matrices in $GL({\mathbb{Z}},3)$). Now, $\mathcal{KT}=K(\pi,1)$ since it is a nilmanifold and $G(\mathcal{KT})={\mathbb{Z}}^{2}$ since the center of $\mathcal{H}_{3}$ is ${\mathbb{Z}}$. Therefore, being an Eilenberg–Mac Lane space is not sufficient to obtain $\operatorname{{\mathsf{rank}}}(G(X))=\operatorname{{\mathsf{dcat}}}(X)$. Also notice that $b_{1}(\mathcal{KT})=3\not=\operatorname{{\mathsf{rank}}}(G(\mathcal{KT}))$ since nilmanifolds are not flat, just almost flat.
@@
Suppose $M=L\times Y$ for a closed manifold $Y$ such that $\pi_{1}(Y)$ is finite with a trivial center and a closed manifold $L$ with an H-space structure (for example, a closed Lie Group). Then
$\operatorname{{\mathsf{dcat}}}(M)>\operatorname{{\mathsf{rank}}}(\pi_{1}(L)).$
@@
First, note that such a $Y$ always exists since any finitely presented group can be realized as the fundamental group of a closed $4$-manifold. For such an $L$, we have $G(L)=\pi_{1}(L)$. Because $\mathcal{Z}\pi_{1}(Y)=0$, $G(M)=G(L)=\pi_{1}(L)$ has finite index in $\pi_{1}(M)$. By Theorem 4.2 (3), since $M$ is not a torus, we must have $\operatorname{{\mathsf{rank}}}(\pi_{1}(L))=\operatorname{{\mathsf{rank}}}(G(M))<\operatorname{{\mathsf{dcat}}}(M)$. ∎
The shortness of the proof of Corollary 4.4 belies the fact that there does not seem to be another way of seeing $\operatorname{{\mathsf{dcat}}}(M)>\operatorname{{\mathsf{rank}}}(\pi_{1}(L))$. Note that, in particular, if $L=T^{k}$ above, we get $\operatorname{{\mathsf{dcat}}}(M)>k$. Moreover, the closed Lie group $L$ can be chosen such that $\pi_{1}(L)$ is any given finitely generated abelian group using tori and the fact that the projective special unitary group has $\pi_{1}(PSU(n))={\mathbb{Z}}/n{\mathbb{Z}}$.
@@
If a torus $T^{k}$ acts on a closed $n$-manifold $M$ such that the map ${\mathcal{O}}_{*}:\pi_{1}(T^{k})={\mathbb{Z}}^{k}\to\pi_{1}(M)$ induced by the orbit map ${\mathcal{O}}$ is a monomorphism, the action is called an injective toral action. These were studied extensively by Conner–Raymond in [CR71]. By (3) of Section 2.E, any action has image $\text{Im}(\mathcal{O}_{*})\subset G(M)$ So, $G(M)$ contains a copy of ${\mathbb{Z}}^{k}$. Therefore, we have the situation of Theorem 4.2 and there is an associated covering $M^{\prime}\simeq T^{k}\times N$ with $N$ simply connected. It was shown in [Op02b] that if $M$ is a closed $n$-manifold with an injective toral action of the torus $T^{k}$, then by the arguments as above, $\operatorname{{\mathsf{cat}}}(M)\geq k+\operatorname{{\mathsf{cup}}}(N)$, and the equality $\operatorname{{\mathsf{cat}}}(M)=k$ holds if and only if $k=n$ and $M\simeq T^{n}$. Any flat manifold $M$ has an injective toral action by $T^{k}$, where $k=b_{1}(M)$. For $\operatorname{{\mathsf{dcat}}}$, however, we only have conclusions as in Theorem 4.2 for injective total actions.
@@
where ${\mathbb{Z}}/s_{i}{\mathbb{Z}}\subset S^{1}\subset S^{3}$ gives the factor space $S^{3}/({\mathbb{Z}}/s_{i}{\mathbb{Z}})$. It is known that the Gottlieb group of a factor space is the center of the group, so we see that $G(S^{3}/({\mathbb{Z}}/s_{i}{\mathbb{Z}}))={\mathbb{Z}}/s_{i}{\mathbb{Z}}$ itself. By Property (4) in Section 2.E, we then have $G(M)=A$. But now we notice the following two things about $M$.
First, since each quotient $S^{3}/({\mathbb{Z}}/s_{i}{\mathbb{Z}})$ is a closed orientable $3$-manifold, we have $H^{3}(S^{3}/({\mathbb{Z}}/s_{i}{\mathbb{Z}});{\mathbb{Q}})={\mathbb{Q}}$ and $H^{k}(S^{3}/({\mathbb{Z}}/s_{i}{\mathbb{Z}});{\mathbb{Q}})=0$ in degrees $1$ and $2$ (this follows from the transfer in Lemma 3.5). Thus, $\operatorname{{\mathsf{cup}}}(M)=k+n$. Therefore,
@@
$H\to L\to L/H\to BH\to BL,$
so that the connecting homomorphism of the fibration $L\to BH\to BL$ may be taken to be $\partial_{*}\colon\pi_{1}(L)\to\pi_{1}(L/H)$. This implies that $\text{Im}(\partial_{*})\subset G(L/H)$. Since $H$ is connected, $\partial_{*}$ is an epimorphism. Thus, $G(L/H)=\pi_{1}(L/H)$.

### v2

@@ 4. Extensions for distributional category
In this section, we will see that $\operatorname{{\mathsf{dcat}}}$ allows us to improve the results of Section 3. As motivation, recall that a compact Riemannian manifold $M^{n}$ is flat if the sectional curvature at each point is zero. It is known that $M^{n}$ is flat if and only if $M\cong{\mathbb{R}}^{n}/G$, where $G$ is a Bieberbach group (see below) acting freely on ${\mathbb{R}}^{n}$ and $\cong$ denotes “is isometric to”. Hence, $M$ is aspherical with fundamental group $\pi_{1}(M)\cong G$. Bieberbach’s Theorem characterizes a Bieberbach group $G$ as one having a maximal normal free abelian subgroup $N\cong{\mathbb{Z}}^{k}$ such that $G/N=Q$ is finite. In [Vas70], Vasquez removed the maximal condition by showing that any free abelian $N$ with $G/N$ finite has a normal maximal free abelian centralizer of the same rank. Furthermore, using the fact that a finite index subgroup contains a normal finite index subgroup (namely the normal core), it can be shown that $G$ is the fundamental group of a compact flat Riemannian $n$-manifold if and only if
@@
Hence, if $M$ is flat, then $\pi_{1}(M)=G$ obeys these conditions. The converse was shown in [FH83]; that is, a closed (connected) manifold $M^{n}$ ($n\not=3,4$) [^1] For the case $n=3$, see [Wol67], and for the case $n=4$, see [Lam07]. supports a flat Riemannian structure if and only if $M^{n}$ is aspherical (with $\operatorname{{\mathsf{cd}}}(\pi_{1}(M))=n$) and $\pi_{1}(M)$ contains an abelian subgroup of finite index. From Section 2.E, we see that a flat manifold $M$ has the Gottlieb group given by the center of the fundamental group, i.e., $G(M)=\mathcal{Z}\pi_{1}(M)$. Moreover, by Theorems 2.4 and 3.7, we have
@@
In [HS86], it was shown that $\operatorname{{\mathsf{rank}}}(\mathcal{Z}\pi_{1}(M))=b_{1}(\pi_{1}(M))=b_{1}(M)$, so we see that we also have $b_{1}(M)\leq\operatorname{{\mathsf{dcat}}}(M)$, an analog to Theorem 3.3. In this case, the condition $b_{1}(M)=\operatorname{{\mathsf{cat}}}(M)=\operatorname{{\mathsf{dcat}}}(M)=\dim(M)=n$ then says that $\pi_{1}(M)$ contains a central free abelian subgroup ${\mathbb{Z}}^{n}$. This can only happen if $\pi_{1}(M)$ is itself free abelian of rank $n$ (essentially by the same argument as in the proof of Theorem 4.2 (3) given below). Hence, $M\cong T^{n}$.
Since flat manifolds have non-negative Ricci curvature, they provide non-trivial examples of such manifolds with non-trivial Gottlieb groups, where we can also determine $\operatorname{{\mathsf{dcat}}}$. Also, in [DHS09] many flat manifolds were described that are Kähler (hence symplectic), so we have such non-trivial examples of $c$-symplectic nature too. Finally, we note that many of these flat Kähler manifolds have non-trivial centers (determined by their non-zero first Betti numbers), so they have non-trivial Gottlieb groups as well.
@@
Let $M$ be a closed $n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is infinite. If $M^{\prime}\cong T^{r}\times W$ is a Cheeger–Gromoll splitting of $M$, then
@@
If $M$ is a torus, then $b_{1}(M)=\operatorname{{\mathsf{cat}}}(M)=\operatorname{{\mathsf{dcat}}}(M)$ by Theorems 3.3 and 2.4. Conversely, if $b_{1}(M)=\operatorname{{\mathsf{dcat}}}(M)$, then the above inequalities give $\operatorname{{\mathsf{cup}}}(W)=0$. Since $W$ is closed and simply connected (hence orientable), this implies $W\simeq\ast$. Hence, $M^{\prime}\cong T^{n}$. This means $M$ is a closed aspherical $n$-manifold. Therefore, $b_{1}(M)=\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)=n$ by Theorem 2.4. But this is the situation of Bochner’s theorem, so we obtain $M\cong T^{n}$. ∎
@@
Recall that in geometric group theory (see [BB97]), a discrete group $\Gamma$ is said to be in $FH(R)$ (written $\Gamma\in FH(R)$) for a commutative ring $R$ if $\Gamma$ acts freely, properly discontinuously, cellularly, and cocompactly on an $R$-acyclic space (i.e., a CW complex $Y$ with $H^{j}(Y;R)=0$ for $j>0$). This is one step in the ladder of finiteness conditions for a group (also see [Bro82]).
@@
Let $p:X^{\prime}\to X$ be a regular covering map corresponding to $A\subset\pi_{1}(X)$, so that $\pi_{1}(X^{\prime})\cong A$. In fact, by Property (2) in Section 2.E, we have $\pi_{1}(X^{\prime})=G(X^{\prime})$ since $p_{*}(A)\subset G(X)$. Because $A$ is free abelian, we have $A\cong{\mathbb{Z}}^{k}$ for some $k$ so that the Hurewicz rank of $A$ is $k$. Thus, by Theorem 2.5, we have $X^{\prime}\simeq T^{k}\times Y$, where $Y$ is simply connected. Clearly, Theorem 2.2 gives us the following series of inequalities (where recall that $\operatorname{{\mathsf{cup}}}$ denotes the rational cup-length):
@@
Note that in the previous argument, we only needed $X$ to be compact to fit the definition of $FH({\mathbb{Q}})$. The fact that $Y$, and hence $\widetilde{X}$, is $\mathbb{Q}$}-acyclic does not depend on the compactness of $X$. Now, if $X$ is a nilpotent space (i.e., $\pi_{1}(X)$ is nilpotent and acts nilpotently on the homology of the universal cover), we can rationalize it to get $X_{\mathbb{Q}}$ with the property that $\pi_{1}(X_{\mathbb{Q}})$ is the Malcev $\mathbb{Q}$}-completion $\pi_{1}(X)_{\mathbb{Q}}$ of the nilpotent group $\pi_{1}(X)$, and $\pi_{j}(X_{\mathbb{Q}})\cong\pi_{j}(X)\otimes{\mathbb{Q}}\cong\pi_{j}(\widetilde{X})\otimes{\mathbb{Q}}\cong\{1\}$ for each $j\geq 2$ because $\widetilde{X}$ is simply connected and $\mathbb{Q}$}-acyclic. Therefore, $X_{\mathbb{Q}}\cong K(\pi_{1}(X_{\mathbb{Q}}),1)\cong K(\pi_{1}(X)_{{\mathbb{Q}}},1)$ is an Eilenberg–Mac Lane space and $X$ itself is rationally homotopy equivalent to $K(\pi_{1}(X),1)$. This is often written as $X\simeq_{\mathbb{Q}}K(\pi_{1}(X),1)$.
Now, suppose that $X$ is a closed $n$-manifold and $G(X)$ has finite index in $\pi_{1}(X)$. Then we see that $A\cong{\mathbb{Z}}^{k}$ (as above) also has a finite index in $\pi_{1}(X)$. Therefore, $X^{\prime}\simeq T^{k}\times Y\to X$ is a finite covering and $X^{\prime}$ is then a closed $n$-manifold as well. Thus, $Y$ is also a closed manifold of dimension nkn-k. Just as before, since $\operatorname{{\mathsf{rank}}}(G(X))=k=\operatorname{{\mathsf{dcat}}}(X)$, we have $\operatorname{{\mathsf{cup}}}(Y)=0$. Hence, $Y$ is $\mathbb{Q}$}-acyclic and simply connected. This means that $X^{\prime}$ is an orientable closed $n$-manifold. So, we must have $H_{n}(X^{\prime};{\mathbb{Q}})\cong{\mathbb{Q}}$. But since $Y$ is $\mathbb{Q}$}-acyclic, this can only happen if $k=n$, $Y\simeq*$, and $X^{\prime}\simeq T^{n}$. This then means that $X\simeq K(\pi_{1}(X),1)$, so that $\pi_{1}(X)$ is torsion-free. But then we have a short exact sequence
@@
where ${\mathbb{Z}}^{n}$ is central in $\pi_{1}(X)$ and $F$ is finite. Consider the associated $5$-term exact sequence (see [Bro82, Corollary VII.6.4])
@@
Because ${\mathbb{Z}}^{n}$ is central, the middle term reduces to ${\mathbb{Z}}^{n}$ itself. The quotient $F$ is finite, so $H_{2}(F)$ is finite as well and therefore, the image of $H_{2}(F)\to{\mathbb{Z}}^{n}$ is zero and ${\mathbb{Z}}^{n}\to H_{1}(\pi_{1}(X))$ is an injection. This translates to an injection ${\mathbb{Z}}^{n}\to\pi_{1}(X)/[\pi_{1}(X),\pi_{1}(X)]$ so that ${\mathbb{Z}}^{n}\cap[\pi_{1}(X),\pi_{1}(X)]=\{1\}$. But then we have an injection $\pi_{1}(X),\pi_{1}(X)]\to\pi_{1}(X)/{\mathbb{Z}}^{n}$, where $\pi_{1}(X)/{\mathbb{Z}}^{n}$ is finite by our hypothesis. Since as a subgroup of $\pi_{1}(X)$, $\pi_{1}(X),\pi_{1}(X)]$ is torsion-free, we must have $\pi_{1}(X),\pi_{1}(X)]=\{1\}$. Therefore, $\pi_{1}(X)$ is a finitely generated torsion-free abelian group. This implies that $\pi_{1}(X)\cong{\mathbb{Z}}^{n}$ and $X\simeq T^{n}$. ∎
@@
If $M$ is a closed nilmanifold, then $\operatorname{{\mathsf{dcat}}}(M)=\dim(M)$ by Theorem 2.4. Note that $\operatorname{{\mathsf{rank}}}(G(M))\leq\operatorname{{\mathsf{rank}}}(\mathcal{Z}\pi_{1}(M))<\dim(M)$, except when $M$ is a torus, see [OS12, Appendix B]. Therefore, being a compact Eilenberg–Mac Lane space is not sufficient to obtain $\operatorname{{\mathsf{rank}}}(G(X))=\operatorname{{\mathsf{dcat}}}(X)$.
@@
Suppose $M=L\times Y$ for a closed manifold $Y$ such that $\pi_{1}(Y)$ is finite with a trivial center and a closed manifold $L$ with an H-space structure (for example, a closed Lie Group). Then $\operatorname{{\mathsf{dcat}}}(M)>\operatorname{{\mathsf{rank}}}(\pi_{1}(L))$.
@@
First, note that such a $Y$ always exists since any finitely presented group can be realized as the fundamental group of a closed $4$-manifold. For such an $L$, we have $G(L)=\pi_{1}(L)$. Because $\mathcal{Z}\pi_{1}(Y)=\{1\}$, $G(M)=G(L)=\pi_{1}(L)$ has finite index in $\pi_{1}(M)$. By Theorem 4.2 (3), since $M$ is not a torus, we must have $\operatorname{{\mathsf{rank}}}(\pi_{1}(L))=\operatorname{{\mathsf{rank}}}(G(M))<\operatorname{{\mathsf{dcat}}}(M)$. ∎
The shortness of the proof of Corollary 4.4 belies the fact that there does not seem to be another way of seeing $\operatorname{{\mathsf{dcat}}}(M)>\operatorname{{\mathsf{rank}}}(\pi_{1}(L))$. Note that, in particular, if $L=T^{k}$ above, we get $\operatorname{{\mathsf{dcat}}}(M)>k$. Moreover, the closed Lie group $L$ can be chosen such that $\pi_{1}(L)$ is any given finitely generated abelian group using tori and the fact that the projective special unitary group has $\pi_{1}(PSU(n))\cong{\mathbb{Z}}/n{\mathbb{Z}}$.
@@
If a torus $T^{k}$ acts on a closed $n$-manifold $M$ such that the map ${\mathcal{O}}_{*}:\pi_{1}(T^{k})\cong{\mathbb{Z}}^{k}\to\pi_{1}(M)$ induced by the orbit map $\mathcal{O}$} is a monomorphism, the action is called an injective toral action. These were studied extensively in [CR71]. By (3) of Section 2.E, any such action has $\text{Im}(\mathcal{O}_{*})\subset G(M)$. Thus, $G(M)$ contains a copy of ${\mathbb{Z}}^{k}$. So, we have the situation of Theorem 4.2 and there is an associated covering $M^{\prime}\simeq T^{k}\times N$ with $N$ simply connected. It was shown in [Opr02b] that $\operatorname{{\mathsf{cat}}}(M)\geq k+\operatorname{{\mathsf{cup}}}(N)$, and $\operatorname{{\mathsf{cat}}}(M)=k$ holds if and only if $k=n$ and $M\simeq T^{n}$. Any flat manifold $M$ has an injective toral action by $T^{k}$, where $k=b_{1}(M)$. For $\operatorname{{\mathsf{dcat}}}$, however, we only have conclusions as in Theorem 4.2 for injective total actions.
@@
where ${\mathbb{Z}}/s_{i}{\mathbb{Z}}\subset S^{1}\subset S^{3}$ gives the factor space $S^{3}/({\mathbb{Z}}/s_{i}{\mathbb{Z}})$. For this type of factor space, the Gottlieb group is the center of ${\mathbb{Z}}/s_{i}{\mathbb{Z}}$ by [Lan72][Opr91], so we see that $G(S^{3}/({\mathbb{Z}}/s_{i}{\mathbb{Z}}))\cong{\mathbb{Z}}/s_{i}{\mathbb{Z}}$ itself. By Property (4) in Section 2.E, we then have $G(M)\cong A$. But now we notice the following two things about $M$.
First, since each quotient $S^{3}/({\mathbb{Z}}/s_{i}{\mathbb{Z}})$ is a closed orientable $3$-manifold, we have $H^{3}(S^{3}/({\mathbb{Z}}/s_{i}{\mathbb{Z}});{\mathbb{Q}})\cong{\mathbb{Q}}$ and $H^{k}(S^{3}/({\mathbb{Z}}/s_{i}{\mathbb{Z}});{\mathbb{Q}})=\{1\}$ in degrees $1$ and $2$ (this follows from the transfer in Remark 3.4). Thus, $\operatorname{{\mathsf{cup}}}(M)=k+n$. Therefore, we have that
@@
$\Omega(BH)\simeq H\to\Omega(BL)\simeq L\to L/H\to BH\to BL,$
so that the connecting homomorphism of the fibration $L/H\to BH\to BL$ may be taken to be $\partial_{*}\colon\pi_{1}(L)\cong\pi_{1}(\Omega(BL))\cong\pi_{2}(BL)\to\pi_{1}(L/H)$. This implies that $\text{Im}(\partial_{*})\subset G(L/H)$ — see Property (5) in Section 2.E. Since $H$ is connected, $\partial_{*}$ is an epimorphism. Thus, $G(L/H)\cong\pi_{1}(L/H)$.